how to solve (6x^3+2x^2-48x-36) =0
and help me on how to solve a 1) Linear equation
(2) Quadratic equation
(3) Polynomial equation
(4) Trigonometric equation
(5) Radical equation
(6) Exponential equation

Grade:12th pass

1 Answers

Saurabh Koranglekar
5 years ago
A+B+C = -b/a
AB+BC+AC=c/a
ABC=-d/a

even if you get one factor you can divide the polynomial by standard division practice

factorise quotient and you will get 3 factors